OpenStudy (maleena_1):

Write the resultant of the two vectors as an ordered pair? <5.-2> & <0,0>

OpenStudy (anonymous):

make a guess, i bet you are right hint: the zero vector acts just like 0

OpenStudy (maleena_1):

what?

OpenStudy (anonymous):

it is just the first one \(<0,0>\) doesnt change anything

OpenStudy (maleena_1):

ok I need how to do it????

OpenStudy (anonymous):

put them end to end, but the zero vectors is just at the origin so there is nothing to do

OpenStudy (maleena_1):

that doesn't help

zepdrix (zepdrix):

\[\large\rm \left<5,-2\right>+\left<0,0\right>\quad=\quad \left<5+0,-2+0\right>\quad=\quad \left<5,-2\right>\]